A cyclist rides 3 km west and then turns around and rides 2 km east.





(a) Draw the FBD for this motion (Assume east to be positive and west to be





negative).





(b) What is her displacement?





(c) What distance does she ride?





(d) What is the magnitude of her displacement?

1
Expert's answer
2022-01-12T08:29:10-0500

"d=-3+2=-1~km,"

"l=3+2=5~km,"

"s=|d|=1~km."